Which of the following indicates the existence of
Answer #1

32)

given

there are strong intermolecular forces

So

the molecules are not free to move , as they are bound to each other

As a result

the molecules cannot any pressure

So the vapor pressure is low

the answer is

A) a very low vapor pressure

33)

B) CH3CH2CH2F has the lowest normal boiling point

34)

C) CH3NH2 has the high normal boiling point

35)

E) SiH4 has the weakest intermolucular forces


36)

we know that

higher the vapor pressure , lower the boiling point

in the given data

diethyl ether has the lowest boiling point

So

C) diethyl ether has the highest vapor pressure at room temperature

37)

we know that

PV/T = constant

so

at two different conditions

P1V1/T1 =P2V2/T2

so

730 x 2 / 296 = P2 x 2 / 356

P2 = 877.97 mm Hg

so

the pressure at 83 C is 878 mm Hg

so

the answer is B) 878 mm Hg

38)

we know that

PV/T = constant

So

P1V1 / T1 = P2V2 / T2

so

1030 x 166 / 298 = P2 x 730 / 298

P2 = 234.2

so

the final pressure is 234.2 mm Hg

so

the answer is E) 234 mm Hg